OpenStudy (anonymous):

@simpleish please answer 6, 7, 10

OpenStudy (anonymous):

6. Solve for the Angle of the incline first by using: Tanθ = v^2/rg. Make sure you convery km/h to m/s for v^2. r = radius. and g = 9.8 Then, use F = mv^2/r After a ton of substitution and simplifying, you should get: sqrt(rg(Tanθ -Ustatic)/ (1 + (Ustatic*Tanθ)))

OpenStudy (anonymous):

7. Same equation for number six, but instead, signs are flipped. Soo sqrt(rg(Tanθ + Ustatic)/ (1 - (Ustatic*Tanθ)))

OpenStudy (anonymous):

10. Fnet = mv^2/r

OpenStudy (anonymous):

and you get (6) Tan(theta) = v^2 / rg from Nsin(theta) = mv^2/r and using Ncos(theta) = mg to substitute for the N in the first one

OpenStudy (anonymous):

pst simpleish, nina's here!

OpenStudy (anonymous):